OpenStudy (anonymous):

What are the conditions needed in a binomial experiment? Give some examples of binomial probabilities.

OpenStudy (campbell_st):

there needs to be only 2 possible outcomes and there also needs to be repeated trials. e.g rolling a bias die where the probability of rolling a 6 is p = 1/3. If the die is rolled 4 times what is the probability of no '6's appearing so not 6 q = 2/3 using binomial expansion of (p + q)^4 = p^4 + p^3q + p^2q^2 + pq^3 + q^4 the answer is (2/3)^4 = 16/81 in general its just